CHIRAL SELF-RECOGNITION: DIRECT ROTATIONAL SPECTROSCOPIC DETECTION OF 2-FLUOROETHANOL DIMERS AND TRIMERS IN THE GAS PHASE

\underline{XUNCHEN LIU}, NICOLE BORHO, YUNJIE XU, Department of Chemistry, University of Alberta, Edmonton, Alberta, Canada, T6G 2G2.

The 2-fluoroethanol (FE) monomer has five conformers and exists predominately in the all-gauch conformation which is stabilized by an intra-molecular O-H^{...}F bond. When interacting with another molecule through hydrogen bonding, FE exhibits transient chirality.^{a} In this study, we examine the competition between the intra- and inter-molecular hydrogen bonds when two FE form a dimer using rotational spectroscopy and ab initio calculations. Three pairs of the homoand heterochiral binary conformations are predicted to be local minima at the MP2/aug-ccpVTZ//MP2/6-311++G(d,p) level of theory. Jet-cooled rotational spectra of the four out of six binary conformers were measured and unambiguously assigned for the first time. All four observed binary conformers are in the compact conformation, where two binding subunits have large contact area. The four conformers have very similar binding energies. These observations are in agreement with the previous jet-cooled low resolution infrared study by Suhm and co-workers.^{b} The FE trimers, on the other hand, show strong preference for the homochiral trimers where FEs can form cooperative O-H^{...}O and C-H^{...}F rings. Ab initio calculation shows that the C_{3} symmetric trimer of the all-gauch conformation is the most stable form. The comparison between the binary and ternary FE systems will be discussed.

aa) Paul Hearn, Rachel Cobley, Brian Howard, J. Chem. Phys. 2005, 123, 134324; b) Nicole Borho, Yunjie Xu, Angew. Chem.

Int. Ed. 2007, 46, 2276-2279.

ba) Tina Scharge, Corinna Emmeluth, Thomas Haber, Martin A. Suhm, J. Mol. Struct. 2006, 786 86-95. b) Nicole Borho, Martin A. Suhm, Org. Biomol. Chem., 2003, 1, 4351-4358.